Table below contains the information on decomposition of ammonia. Using that table calculate the initial and...

  1. Table below contains the information on decomposition of ammonia. Using that table calculate the initial and average rate of reaction.

                             

Time, min

NH3, M

0

6

1

3

2

1.5

3

0.75

4

0.375

5

0.1875


Show your work.
Answer #1

Solution:

From the table,

Part A) When, t = 0 min, then [NH3] = 6 M

At ,t = 1 min, [NH3] = 3 M

Thus,

Initial rate = - dX/dt = - (3 M - 6 M) / 1 - 0 min

Initial rate = 3.0 M /min

Part B)

Average rate = - dX/dt

dX = final concentration - initial concentration

dX = 0.1875 M - 6.0 M = -5.8125 M

dt = 5 -0 = 5 min

Therefore,

Average rate = - (- 5.8125) M / 5 min

= 1.1625 M/min